OpenStudy (anonymous):

A tree casts a shadow measuring 555 ft. At the same time, a five ft. tall person casts a shadow of 7.5 feet. How tall is the tree? All I know is that the person's shadow is 1.5 times longer than the person itself.

OpenStudy (anonymous):

Good! That means that the tree's shadow is also 1.5 times longer than the tree itself... Can you figure out the tree height now? You could also solve this problem by setting up a proportion and cross-multiplying.

OpenStudy (anonymous):

And the proportion setup would be 7.5 over 555 is equal to x over 5 ???

OpenStudy (anonymous):

Almost. If you do one fraction for the shadows and one fraction for actual heights, you have to make sure you are putting the same "thing" in the numerator/denominator of each fraction. Meaning, you need to do: \[ \frac{ man shadow }{ tree shadow } = \frac{ man height }{ tree height }\] The proportion YOU set up is: \[\frac{ man shadow }{ tree shadow } = \frac{ tree height }{ man height }\] but you're not allowed to do that -- you're not comparing the same things in the same parts of your fraction. Do you see the difference?

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

You're welcome! :)